George Robbins considers himself an aggressive investor.​ He's thinking about investing in some foreign securities and is looking at stocks in​

(1) Bayer​ AG, the big German chemical and​ health-care firm, and​

(2) Swisscom​ AG, the Swiss telecommunications company. Bayer​ AG, which trades on the Frankfurt​ Exchange, is currently priced at 57.72 euros ​(euro ​) per share. It pays annual dividends of 1.52 euro per share.

Robbins expects the stock to climb to 65.57 euro per share over the next 12 months. The current exchange rate is 0.9092 euro ​/U.S. ​$, but​ that's expected to rise to 0.9888 euro ​/U.S. ​$.

The other​ company, Swisscom, trades on the Zurich Exchange and is currently priced at 71.33 Swiss francs​ (Sf) per share. The stock pays annual dividends of 1.48 Sf per share. Its share price is expected to go up to 75.16 Sf within a year. At current exchange​ rates, 1 Sf is worth $ 0.7616 ​U.S., but​ that's expected to go to $ 0.8554 by the end of the​ 1-year holding period.

a. Ignoring the currency effect​, which of the two stocks promises the higher total return​ (in its local​ currency)? Based on this​ information, which of the two stocks looks like the better​ investment?

b. Which of the two stocks has the better total return in U.S. dollars​? Did currency exchange rates affect their returns in any​ way? Do you still want to stick with the same stock you selected in part a​? Explain.
Answer #1

a. Ignoring the currency effect​, which of the two stocks promises the higher total return​ (in its local​ currency)? Based on this​ information, which of the two stocks looks like the better​ investment?
Bayer=(65.57+1.52)/57.72-1=16.23%

Swisscom=(75.16+1.48)/71.33-1=7.44%

Bayer looks like better investment

b. Which of the two stocks has the better total return in U.S. dollars​? Did currency exchange rates affect their returns in any​ way? Do you still want to stick with the same stock you selected in part a​? Explain.
Bayer=((65.57+1.52)/0.9888)/(57.72/0.9092)-1=6.88%

Swisscom=(75.16+1.48)*0.8854/(71.33*0.7616)-1=24.91%

Swisscom looks like better investment